OpenStudy (anonymous):

Simplify. Write as products of powers with no variables in the denominator. 6/x^-4 + 5/x^-2

OpenStudy (anonymous):

6/x^-4 + 5/x^-2 1 = 6x^4 + 5x^2 (because a^m = ----- ) a^-m = x^2[6x^2 + 5]